# Don’t look back: Ordering and defect cloaking in non-reciprocal lattice XY models

### v1

@@ Acknowledgements.
PP acknowledges discussions with Sumantra Sarkar, Navdeep Rana, and Vikash Pandey, and SR and AM with Demian Levis and Sarah Loos. AM acknowledges the support of ANR through the grant PSAM. AM and SR thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the program “Anti-diffusive dynamics: from sub-cellular to astrophysical scales”. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics. AM acknowledges a TALENT fellowship awarded by CY Cergy Paris Université, and SR a JC Bose Fellowship of the ANRF, India.
@@ Appendix A Expressions for the angle field equation for triangular lattices
Appendix B Simulation details
@@
Appendix C Details of supplementary videos

### v2

@@ Acknowledgements.
PP acknowledges discussions with Sumantra Sarkar, Navdeep Rana, and Vikash Pandey, and SR and AM with Demian Levis and Sarah Loos. AM acknowledges the support of ANR through the grant PSAM. AM and SR thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the program “Anti-diffusive dynamics: from sub-cellular to astrophysical scales” (EPSRC grant no. EP/R014604/1), and SR additionally for a Rothschild Distinguished Visiting Fellowship. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics. AM acknowledges a TALENT fellowship awarded by CY Cergy Paris Université, and SR a JC Bose Fellowship of the ANRF, India.
@@ Appendix A Expressions for the angle field equation for triangular lattices
Appendix B Additional measurements of the properties of the ordered phase in our model and in the model of [17]
Figure 9: Statistical properties of the ordered phase on a triangular lattice. (a) Two-point correlation function $C(r)$ for various system sizes $L$, with distance $r$ in units of $L/2\pi$, approaches a constant nonzero value for large $r$, indicating long-range order. Averaging is carried out over five independent noise realizations. Inset shows magnetization $m(L)\equiv\sqrt{C(r=L)}$ as a function of $L^{-1}$. (b) Measured structure factors $S_{\perp}$ as functions of $k_{\parallel}$ and $k_{\perp}$ for $L=512$. The analysis is performed at reduced temperature $T=0.2$ and nonreciprocity strength $\mathcal{A}=1.5$, with time averaging over four independent noise realizations. (c) Measured structure factors $S_{\perp}$ as functions of $k_{\parallel}$ and $k_{\perp}$ for $L=512$, with time averaging over four independent noise realizations for $\mathcal{A}=3.5$. The flattening out of the structure factor at small wavenumbers was not very clear for $\mathcal{A}=2.5$ in the triangular lattice, in the wavenumber regime we examine; therefore, we display the result for $\mathcal{A}=3.5$ where we observe that the scaling of the structure factor becomes independent of $k_{\parallel}$ at small $k_{\parallel}$.
Figure 10: Measured structure factors $S_{\perp}$ as functions of $k_{\parallel}$ and $k_{\perp}$ for $L=512$ for (a) square and (b) triangular lattice for the model in [17]. The analysis is performed at reduced temperature $T=0.2$ and $\sigma=2.5$ with time-averaging over five noise realisations. The flattening of $S_{\perp}$ at small wavenumbers clearly implies a nonzero relaxation rate of angular fluctuations in this limit due to the presence of a term $\propto\theta$ in this case.
Appendix C Simulation details
@@
Appendix D Details of supplementary videos